Grade 12Algebra

can you please explain the important concepts of inverse trignometric functions and also with their formulaes?

revathy t.nair
14 Years agoGrade 12

1 Answer

Aman  Bansal
14 Years ago

Daar Student,

The functions sin-1x, cos-1x, tan-1x, cot-1x, cosec-1x and sec-1x are called inverse circular or inverse trigonometric functions. sin-1x should not be confused with 

(sin x)-1 which is equal to .

Each of the inverse circular functions is multivalued. To make each inverse circular function single valued we define principal values as follows. If x is positive, the 

principal values of all the inverse circular functions lie between 0 and . If x is negative, the principal values of sin-1x, cosec-1x and tan-1x lie betweenand 0, and 

those of cos-1x, sec-1x and cot-1x lie between  and . From now onwards we take only the principalvalues.

Hence sin = x   = sin-1x where   (range) and x  [-1, 1] (domain).

This ensures that the function  = sin-1x is one-one and onto.

Similarly cosec = x   = cosec-1x where   and x  (-, -1]  [1, ).

cos = x   = cos-1x where   [0, ] and x  [-1, 1].

sec = x   = sec-1x where    and x  (-, -1]  [1, ).

tan = x   = tan-1x where   and x  (-, ).

cot = x   = cot-1x where   (0, ) and x  (-, ).


The domains and ranges of inverse trigonometric (or inverse circular) functions are:

Function Domain Range
sin-1 x [- 1, 1]  
cos-1 x [- 1, 1] [0, ]
tan-1 x R  
cot-1 x R (0, )
sec-1 x (- , - 1]  [1, )  
cosec-1 x (- , - 1]  [1, )  


The graphs of  = sin-1x,  = cos-1x and  = tan-1x are:

     

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert